Given\\ (f(x)=3x(x-1)^{5} compute \\ (f\`\`\`(x)\\)


1
Expert's answer
2019-09-09T11:03:24-0400

"f(x)=3x(x-1)^5"

"f'(x)=3(x-1)^5+3x\cdot 5(x-1)^4="

"=3(x-1)^4(x-1+5x)=3(x-1)^4(6x-1)"

"f''(x)=3\cdot 4(x-1)^3(6x-1)+3(x-1)^4\cdot 6="

"=6(x-1)^3(12x-2+3x-3)=30(x-1)^3(3x-1)"

"f'''(x)=30\cdot 3(x-1)^2(3x-1)+30(x-1)^3\cdot 3="

"=90(x-1)^2(3x-1+x-1)=180(x-1)^2(2x-1)"

Answer: "f'''(x)=180(x-1)^2(2x-1)"